OpenStudy (anonymous):

The second problem I am having is: a) solve x^2+7x-2=0 b) find the x intercepts of f(x)=x^2+7x-2 The answer I got for this one was -7+sqrt(5)/2 and -7-sqrt(5)/2. And it told me I was wrong...help.

OpenStudy (mertsj):

Did you use the quadratic formula? a = 1, b = 7, c = -2

OpenStudy (anonymous):

I thought I did.

OpenStudy (mertsj):

\[x=\frac{-7\pm \sqrt{49-4(1)(-2)}}{2}=\frac{-7\pm \sqrt{57}}{2}\]

OpenStudy (anonymous):

x=1/2(-7-(57)^1/2) X=1/2((57)^1/2-7)